OpenStudy (anonymous):

Please Please Help I Will Fan And Give Medal!!!! Its More Than One Answer! What Can You Say About The Y-Values Of The Two Functions f(x)=3x^2-3 And g(x)=2^x-3? A.g(x) has the smallest possible y-value B.f(x) and g(x) have equivalent minimum y-values C.The minimum y-value of g(x) approaches -3 D.f(x) has the smallest possible y-value

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@retirEEd

OpenStudy (retireed):

The second equation g(x) approaches -3, but NEVER gets there. Whereas the first equation f(x) IS -3 at x = 0.

OpenStudy (retireed):

Is it possible to have two correct answers?

OpenStudy (anonymous):

yea its possible @retirEEd

OpenStudy (anonymous):

@Directrix no its well written ?

Directrix (directrix):

I messed up. Sorry. I read g(x)=2^x-3 and g(x) = x^2 - 3. g(x)=2^x-3 is not a parabola.

OpenStudy (retireed):

@lovelyvaal I think Directrix is mis-reading the g(x) equation I see it as ...\[g(x) = 2^{x} - 3\]

OpenStudy (anonymous):

yea he said he read it wrong @retirEEd

OpenStudy (anonymous):

i belive its A and C? @Directrix @retirEEd

OpenStudy (retireed):

Any I think there are two correct answers. What do you think they are?

OpenStudy (retireed):

Yes you beat me to it. Good job!

OpenStudy (anonymous):

A and C? @retirEEd

OpenStudy (retireed):

NO C and D I can't read.

OpenStudy (anonymous):

wait what lol ? A and C? Or C and D? lol Im confused :/? @retirEEd

OpenStudy (retireed):

C g(x) approaches -3 while D is -3, the lowest possible y.

OpenStudy (retireed):

Sorry, I said yes at the wrong answer. C and D are correct.

OpenStudy (anonymous):

Thank you so much ! @retirEEd